HELP NO clue how to do this!
An eagle is circling a field at a height of 250 feet sees a snake below. The eagle dives at an initial speed of 110 feet per second. Determine the time the snake has to escape.

Answer by stanbon(75887)

An eagle is circling a field at a height of 250 feet sees a snake below. The eagle dives at an initial speed of 110 feet per second. Determine the time the snake has to escape.
-----------------------------
Let h(t) be height at time "t".
EQUATION:
h(t) = -16t^2-110t+ 250
When the eagle hits the snake its height is zero.
-16t^2-110t+250 = 0
-2(8t^2+55t-125) = 0
t = [-55 +- sqrt(55^2 - 4*8*-125)]/16

t = [-55 +- sqrt(7025)]/16
t = [-55 +- 83.82]
Positive answer:
t = (-55+83.83)/16
t = 1.8 seconds
=====================
Cheers,
Stan H.